OpenStudy (vijay222ms):

a block slides down a frictionless 30 degre plane in half the time it takes to slide down on an identical rough plane.Then the coefficient of friction between the block and the incline is

OpenStudy (vijay222ms):

@wolf1728 @AlexandervonHumboldt2 @Mimi_x3 @pooja195

OpenStudy (vijay222ms):

@JoannaBlackwelder

OpenStudy (vijay222ms):

@perl @bibby @Nnesha @iambatman @Loser66 @Destinymasha @Kainui

pooja195 (pooja195):

well where do you think you would start?

OpenStudy (vijay222ms):

I have no idea.

OpenStudy (vijay222ms):

I have an another question would you like to try?@pooja195

OpenStudy (vijay222ms):

@jagr2713

jagr2713 (jagr2713):

Yes

OpenStudy (vijay222ms):

Can you help me @jagr2713

jagr2713 (jagr2713):

Sure

OpenStudy (vijay222ms):

Then Please help

jagr2713 (jagr2713):

whats the question

OpenStudy (vijay222ms):

a block slides down a frictionless 30 degree plane in half the time it takes to slide down on an identical rough plane.Then the coefficient of friction between the block and the incline is

jagr2713 (jagr2713):

sorry i dont understand the question but let me get someone else to help you. @iambatman @Destinymasha @Directrix

OpenStudy (vijay222ms):

ok

OpenStudy (anonymous):

I would suggest setting up two differential equations of motion. Have one consider an incline plane with no frictional force and the other consider a frictional force with a coefficient of kinetic friction mu. First consider geometry to figure out what component of gravity will be acting on the block.  We see that the block will slide due to mgsin30. The normal force acting on the block will be mgcos30. I'm not completely confident in this, so hopefully somebody can look this over. But it seems to make sense to me and be pretty straightforward. Write two equations of motion for each incline and solve for position equation. Frictionless:\[ma =mg\sin(30)\]\[a=\frac{g}{2}\]\[v=v_{0}+\frac{g}{2}t\]\[x=x_{0}+v_{0}t+\frac{g}{4}t^{2}\]Assume initial position x0 and initial speed 0 m/s.\[x=x_{0}+\frac{g}{4}t^{2}\] Rough:\[ma=mg\sin(30)-\mu mg\cos(30)\]\[a=\frac{g}{2}-\frac{\sqrt3\mu g}{2}=\frac{g}{2}(1-\sqrt3 \mu)\]\[v=v_{0}+\frac{g}{2}(1-\sqrt3 \mu)t\]\[x=x_{0}+v_{0}t+\frac{g}{4}(1-\sqrt3 \mu)t^{2}\]Assume initial position x0 and initial speed 0 m/s.\[x=x_{0}+\frac{g}{4}(1-\sqrt3 \mu)t^{2}\] Now, we have two equations. We can assume that the block reaches the bottom of the incline (x = 0) for both equations. We can also substitute time 't' into the rough incline equation and 't/2' into the frictionless incline equation. Frictionless:\[0=x_{0}+\frac{g}{4}(\frac{t}{2})^{2}=x_{0}+\frac{g}{16}t^{2}\] Rough:\[0=x_{0}+\frac{g}{4}(1-\sqrt3 \mu)t^{2}\] We equate the two equations and solve for coefficient of friction mu:\[x_{0}+\frac{g}{16}t^{2}=x_{0}+\frac{g}{4}(1-\sqrt3 \mu)t^{2}\]\[\frac{g}{16}t^{2}=\frac{g}{4}(1-\sqrt3 \mu)t^{2}\]\[\frac{1}{4}=1-\sqrt3 \mu\]\[\mu =\frac{\sqrt3}{12}\] So that's what I find for the coefficient. Hope this was the proper approach.